OpenStudy (anonymous):

Which of the following is the equation of a circle whose center is at the origin and whose radius is 5 ? x(squared) + y(squared) =√5 x(squared) + y(squared) = 5 x(squared) + y(squared) = 10 x(squared) + y(squared) = 25

OpenStudy (anonymous):

@peachpi I don't understand how to do this one!

OpenStudy (welshfella):

the equation whose centre is at the origin and whose radius is r is x^2 + y^2 = r^2

OpenStudy (welshfella):

now substitute for r = sqrt5 and you have it

OpenStudy (anonymous):

I got 5

OpenStudy (welshfella):

you got 5 for what - i think you are on the right track

OpenStudy (welshfella):

you need the correct value for r^2

OpenStudy (anonymous):

so then would It be √5?

OpenStudy (welshfella):

√5? is given in the question. its the radius ( r ) of the circle.

OpenStudy (welshfella):

so it √5 is r what is r^2?

OpenStudy (welshfella):

* if

OpenStudy (welshfella):

r^2 = √5 * √5 = ?

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

10?

OpenStudy (welshfella):

^ means 'to the power' or exponential r^2 = r * r

OpenStudy (welshfella):

the square root of some number times the square root of the number = the number ( by definition) so sqrt 3 * sqrt 3 = 3